166 Preliminary Report

Operations Report


The ship departed San Juan at 1709 hr on 20 February 1996. All times reported in this report are in local time (US Eastern Time Zone; UTC-5 hr). The 857-nmi sea voyage to the first way point to begin the seismic survey required approximately 68 hr at 12.5 kt average speed.

A 30-nmi seismic reflection survey was conducted over Sites 1003, 1004, 1005, 1006, and 1008 (Proposed Sites BT-2/F-2, F-3, BT-1A/F-1, BT-20, BT-3, respectively) in 4.7 hr at 6.3 kts. The ship returned to the GPS coordinates for Site 1003.

SITE 1003
(Proposed Site BT-2)

We deployed a Datasonics commandable retrievable beacon at 24°32.746'N, 79°15.642'W on 1812 hr, 23 February 1996. Unless otherwise noted, depths in the operations section within individual site reports refer to meters below rig floor (mbrf). This depth is calculated using the drill pipe measurement (DPM) from the top of the dual elevator stool (DES) on the rig floor. The distance from the dual elevator stool (DES) to sea level varies over time, depending on the ship's draft.

Hole 1003A
An advance hydraulic piston corer/extended core barrel (APC/XCB) bottom hole assembly (BHA) was attached to monel drill collars. Monel drill collars are intended to minimize remagnetization, allowing magnetic orientation of the APC cores. Hole 1003A was spudded at 2215 hr on 23 February. Core 1003A-1H was cut with the drill bit positioned at 490.0 mbrf. Core 1H recovered 6.97 m; therefore, the seafloor was defined to be 492.5 mbrf. The distance between the DES and sea level at Hole 1003A was 11.1 m, making the water depth 481.4 meters below sea level (mbsl).

Adara heat measurements were taken at Cores 3H through 9H. Core orientation using the tensor tool was performed on Cores 4H through 9H. Using the APC, we cored 77.7 m recovering 69.85 m (89.9% recovery). Recovery was reduced by common hard carbonate layers/nodules and occasional sandy carbonate layers in soft carbonate ooze. Core 9H (77.7 mbsf) was a partial stroke with a split core liner and required 60,000 lb overpull to extract the core barrel from the formation. We then switched to the XCB coring method.

XCB cores were cut from Core 10X through 21X (77.7-188.1 mbsf). WSTP temperature measurements were performed after Cores 3X, 6X, 9X, and 22H. Recovery ranged from 0 to 100% (mean = 44%). Cores 17X and 18X had carbonate nodules jammed in the shoe and Cores 20X and 21X had no recovery in a very soft formation, resulting in heavy backflow.

APC Core 22H was taken from 188.1 to 197.6 mbsf with 103% recovery. The APC barrel stuck and could not be extracted from the formation with 60,000 lb overpull. The bit nozzles became plugged with cuttings that flowed back into the pipe, stalling the rotary, and packing off the annulus. Overpull was gradually increased until circulation and rotation were restored. Although several nozzles remained plugged, XCB coring was resumed.

XCB Cores 23X to 27X were cut from 197.6 to 245.7 mbsf with 1.3% recovery. The bit nozzles remained plugged and heavy backflow was observed. Consequently, the hole was terminated and the bit cleared the rig floor at 2200 hr on 24 February.

Hole 1003B
The ship was moved 20 m to the northwest and Hole 1003B was spudded at 0105 hr, 25 February. Core 1003B-1H was taken with the drill bit positioned at 488.0 mbrf. Recovery of the mudline core defined the seafloor at 483.3 mbsl. APC cores were taken from Core 1H to 9H (0 to 79.0 mbsf) with 80.7% recovery. No oriented cores were taken. An Adara temperature tool heat-flow measurement was taken during Core 8H. APC coring was terminated at 79.0 mbsf with 60,000 lb overpull.

XCB cores were cut from Core 10X to 57X (79.0 to 541.6 mbsf) with recovery varying from 0 to 100% (mean=27%). The WSTP was deployed after Core 11X. Recovery was reduced by moderately hard layers that jammed in the shoe. Recovery was low in formation sections that alternated between lithified and unlithified. The pump rate was reduced in soft carbonates to reduce washing and was increased in harder carbonates to reduce jamming. The formation sections became noticeably harder at 34X (310.6 mbsf), 40X (368.2 mbsf), and 52X (483.9 mbsf) with increased drilling time and torque and reduced recovery. Hole 1003B was terminated and the bit cleared the rig floor at 1545 hr, 26 February.

Hole 1003C
The ship was moved 10 m northwest of Hole 1003B. Hole 1003C was spudded using the rotary core barrel (RCB) system at 2020 hr on 26 February. The water depth was assumed to be 483.3 mbsl based on adjacent Hole 1003B. The hole was drilled from 0 to 406.0 mbsf. Cores 1003C-1R to 2R were cut from 406.0 to 425.3 mbsf. After retrieving Cores 1R and 2R, the hole was drilled from 425.3-473.5 mbsf. RCB Cores 3R to 88R were then cut from 473.5 to 1300.0 mbsf. From the original approval depth of 1050 mbsf, Hole 1003C was deepened to 1300 mbsf to reach the base of the Neogene after which coring was terminated. Recovery using the RCB system ranged from 4.1 to 103.0% (mean = 44.7%). At least 21 cores were jammed from fractured core sections wedging in the shoe and liner. A strong petroleum odor was noted in Cores 30R to 41R (732.4-842.0 mbsf), with increasing amounts of heavier hydrocarbons (C2-C6) observed in headspace analyses. C1/C2 ratios decreased from 23 to 8. However, coring continued on a core-to-core basis because the concentration of gas was very low, and there were no rapid changes in gas concentration or C1/C2+ ratios (C1/C2 ratios being already anomalously low from the top of the hole).

During the process of cleaning the hole in preparation for logging, the drill pipe became stuck. For five hours we unsuccessfully attempted to free the pipe using up to the maximum overpull allowed for shallow-water drilling (200,000 lb). The pipe was severed at 1008.5 mbsf with an explosive charge. The severed pipe was pulled to 96 mbsf and we dropped a free-fall funnel (FFF). The vibration-isolation tool was deployed with a television camera and sonar to verify positioning of the FFF and the drill string was pulled out of the hole. A RCB BHA was deployed and the hole was re-entered for logging. However, the drill pipe became stuck again while attempting to condition the hole and we were not able to penetrate deeper than 465 mbsf. We encountered similar conditions while pulling up the drill string suggesting the upper part of the hole was collapsing. Therefore, we judged it imprudent to make any further attempts to log the hole. Hole 1003C was terminated and the bit cleared the rig floor at 1457 hr, 6 March.

Hole 1003D
After evaluating hole stability and logging potential at several sites, we returned to Site 1003 on March 20 to drill a dedicated logging hole. The 14.0 nmi transit from Site 1006 to Site 1003 required 4.25 hr. After positioning the ship on the GPS coordinates (24°32.763'N, 79°15.650'W), a beacon was dropped at 0530, 20 March. A drilling BHA was assembled and run to the seafloor (482.9 mbsl). The hole was drilled to 677.0 mbsf and then conditioned with a short trip from 677.0 to 536.0 mbsf with no overpull or drag. Drilling was resumed from 677.0 to 1052.7 mbsf. A standard wiper trip was performed to 100 mbsf encountering 2.5 m of soft fill, and the bit was released with an MBR. After a circulating a sweep of sepiolite mud, the pipe was pulled to 106 mbsf for logging. Phasor dural induction tool (DIT)/Sonic, Formation MicroScanner (FMS), integrated porosity lithology tool (IPLT), induced gamma ray spectroscopy (GST), and VSP logs were run to 1052.7 mbsf. After plugging the hole with cement and gel mud, the pipe was pulled and secured for transit at 0723 hr on 24 March. Although the beacon indicated release, it did not surface and could not be retrieved.

Site 1004 Operations Report

166 Table of Contents

Publications Home

ODP Home